OpenStudy (anonymous):

A swimming pool has to be drained for maintenance. The pool is shaped like a cylinder with a diameter of 19m and a depth of 2m . If the water is pumped out of the pool at the rate of 15m^3 per hour, how many hours does it take to empty the pool? PLEASE HELP

OpenStudy (anonymous):

Is this problem testing surface area or volume of a cylinder?

OpenStudy (anonymous):

volume vs rate

OpenStudy (anonymous):

yes. Now the volume of a cylinder = (4/3) (pi)(radius)^3

OpenStudy (anonymous):

what will be the volume of the water in this pool?

OpenStudy (anonymous):

Plug the given information into the formula.

OpenStudy (anonymous):

If the diameter is 19, what is the radius of the cylinder?

OpenStudy (anonymous):

Volume of cylinder = (pi)(r^2)(h)... (4/3) (pi)(radius)^3 the same thing?

OpenStudy (anonymous):

sphere?

OpenStudy (anonymous):

Sorry...I stand corrected.. Volume = (4/3)(pi)(radius)^2 (height) (I was thinking we are dealing with a sphere).

OpenStudy (anonymous):

Thanks, @ehuman

OpenStudy (anonymous):

lol I know it's late

OpenStudy (anonymous):

@danitosh Are you working with me on this problem to find the volume?

OpenStudy (anonymous):

I'm just playing referee

OpenStudy (anonymous):

I dont know what happened to this student.

OpenStudy (anonymous):

Guess she'll call when she's ready.

OpenStudy (anonymous):

agree

OpenStudy (anonymous):

Yes! sorry. haha

OpenStudy (anonymous):

the pool is shaped like a cylinder not a sphere

OpenStudy (anonymous):

Volume of a cylinder is (pi)(radius)^2(height) whats the volume of water in this pool? we need radius and height. We are told the diameter is 19. So what is the radius?

OpenStudy (anonymous):

r=9.5

OpenStudy (anonymous):

and the height is 2 meters

OpenStudy (anonymous):

yes. and we know the height is 2. So the volume = (pi)(9.5)^2(2) = what? Use your calculator to compute that.

OpenStudy (anonymous):

567.05

OpenStudy (anonymous):

true.

OpenStudy (anonymous):

water is draining out at 15 cubic meters per hour

OpenStudy (anonymous):

so would it be 5 hrs?

OpenStudy (anonymous):